OpenStudy (anonymous):

Find the x-intercepts using the quadratic formula or FOIL factoring if they exist..

OpenStudy (anonymous):

\[f(x) = -x^2 - 2x + 2\] I get as far as x = -2 + or - the square root of 12 all over 2 when I using the quadratic formula..and now I'm stuck